# The Hydra Map and Numen Formalisms for Collatz-Type Problems

### v1

@@ 1. Document Summary
Section 6 introduces the string formalism (adapted from the theory of automata; see [1]) and uses it to construct the numen, $X_{H}$, of a given Hydra map $H$. The characterizing functional equation of $X_{H}$ is established, and Lemma 1 generalizes the work in [14] to show when and how $X_{H}$ may be extended from a function on $\mathbb{N}_{0}$ to a function on the ring of $p$-adic integers. The section concludes with a statement and sketched proof of the Correspondence Principle, one of the main findings of my earliest work [12, 14]. This relates the integral values taken by $X_{H}$ over $\mathbb{Z}_{p}$ to the periodic and/or divergent points of $H$.
Section 7 begins with a brief review of the $p$-adic Haar probability measure and its use in $p$-adic integration. Several basic $p$-adic integration formulae are introduced, along with the elementary theory of Pontryagin Duality from abstract harmonic analysis and its application to $p$-adic Fourier analysis. Novel content in this section includes Definition 20 on page 20, which gives the $\ell$-adic characteristic function of $X_{H}$, for varying places $\ell$ of the underlying global field $K$. The section also gives the definition of the $\ell$-adic Wiener algebras. These are the spaces of all functions $K_{\ell}\rightarrow\mathbb{C}$ expressible as the inverse Fourier transform of an $L^{1}$ function $K_{\ell}\rightarrow\mathbb{C}$, and provide a natural functional-analytic backdrop for giving a unified treatment of Fourier analysis across the various completions of $K$. Using the Wiener algebras, we can realize the $\ell$-adic characteristic function of $H$ as the Fourier transform of a self-similar Borel probability measure $d\mu_{H,\ell}$. (Equivalently, this measure is the probability distribution associated to the $\ell$-adic valued random variable induced by $X_{H}$.) The section concludes with some useful formulae for working with $X_{H}$ as an $\ell$-adic valued random variable in the case where $\ell$ is a non-trivial non-archimedean place of $K$.
@@ 3. Global Fields, Absolute Values, & Completions
Proposition 4 ($\mu u$ Factorization (Non-Archimedean)).
@@ 5. Hydra Maps
Now we can give the definition of a Hydra map.
Definition 9.
@@
Definition 10.
@@ A $\Lambda$-Hydra (map) on $K$ is a map $H:\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}$ defined by:
Remark 6.
From the perspective of dynamical systems, the integrality property of a Hydra is, by far, its most important property, and plays a critical role in the proof of the Correspondence Principle (Theorem 1, given at the end of Section 6).
@@ 6. The Numen of a Hydra Map
Throughout this section, let $H:\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}$ be a proper $\Lambda$-Hydra map. Following [12, 14], we construct $X_{H}$, the numen of $H$, like so.
Definition 12.
@@
Proposition 9.
$X_{H}\left(\mathbf{i}\right)=X_{H}\left(\mathbf{j}\right)$ for all $\mathbf{i}$ and $\mathbf{j}$ belonging to the same equivalence class in $\Sigma_{p}^{*}/\sim$. Hence, $X_{H}:\Sigma_{p}^{*}/\sim\rightarrow K$ is well-defined. We can then view $X_{H}$ as a function $\mathbb{N}_{0}\rightarrow K$ by using $\textrm{DigSum}_{p}$ to identify $\Sigma_{p}^{*}/\sim$ with $\mathbb{N}_{0}$.
@@
Proposition 10 (Siegel (2022); see [12]).
$X_{H}:\mathbb{N}_{0}\rightarrow K$ is the unique function satisfying the functional equations:
@@
where $r_{j}X_{H}\left(n\right)$ is the image of $X_{H}\left(n\right)\in K$ under $r_{j}\in\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$.
Remark 10.
Proposition 10 only requires $H$ to be proper. Integrality and centeredness are not needed.
@@
In [12], it was shown that, given mild assumptions on the $r_{j}$s, one can use the identification of $\Sigma_{p}^{\infty}/\sim$ with $\mathbb{Z}_{p}$ to interpolate $X_{H}:\mathbb{N}_{0}\rightarrow K$ to a function on $\mathbb{Z}_{p}$. The details for this are given in Lemma 1, below. The essential idea is that by making a judicious choice of absolute values on $K$, one can realize $X_{H}$ as a function (or, in general, as a distribution) out of $\mathbb{Z}_{p}$ whose regularity depends on how close the $r_{j}$s are to behaving as contractions with respect to $K$. The behavior of the $r_{j}$s is captured by their norm with respect to a given absolute value of $K$. The result given below is a slightly modified version of the versions originally stated in [14]. The reader should be aware that Lemma 1 treats $\mathbb{Z}_{p}$ as a measure space equipped with its real-valued Haar probability measure, and treat $K_{\ell}$ as a measure space equipped with its real-valued Haar measure. We also use the concept of a locally constant function. Readers unfamiliar with either of these can find an expository treatment given of them at the start of Section 7.
Lemma 1.
For each non-trivial place $\ell$ of $K$, let:
@@
Now, consider the extension $X_{H}:\mathbb{Z}_{p}\rightarrow K_{\ell}$ defined by the formula
@@
Remark 13.
Like with Proposition 10, Lemma 1 requires only the properness of $H$, as is made clear near the end of the proof.
@@
which shows that
@@
Theorem 1 (Correspondence Principle).
@@ 7. $p$-Adic Fourier Analysis and the $\ell$-adic Characteristic Function of $X_{H}$
Proposition 11 (Decomposition formula).
@@
By Lemma 1, the condition $\rho_{H,\ell}<1$ guarantees that $X_{H}:\mathbb{Z}_{p}\rightarrow K_{\ell}$ is measurable. As such, we can view $X_{H}$ as a $K_{\ell}$-valued random variable whose associated sample space is the Borel $\sigma$-algebra of $\mathbb{Z}_{p}$. Given any measurable subset $A$ of $K_{\ell}$, the probability $\mathbb{F}\left(X_{H}\in A\right)$ would then be the Haar probability measure of the subset of $\mathbb{Z}_{p}$ on which $X_{H}\in A$.
@@
Proof: By (6.25) from Lemma 1, the hypotheses on the $\ell$-adic norms of the $r_{j}$s implies that $X_{H}$ is $\ell$-adically bounded by:
@@
By character orthogonality, using (14), we have:
@@
for all $\mathfrak{w}\in\pi_{\ell}^{-B_{H,\ell}}\mathcal{O}_{K_{\ell}}$. Letting $n\geq B_{H,\ell}$, we can apply Proposition 15 to get:

### v2

@@ 1. Document Summary
Section 6 introduces the string formalism (adapted from the theory of automata; see [1]) and uses it to construct the numen, $X_{H}$, of a given Hydra map $H$. The characterizing functional equation of $X_{H}$ is established, and Lemma 6.1 generalizes the work in [14] to show when and how $X_{H}$ may be extended from a function on $\mathbb{N}_{0}$ to a function on the ring of $p$-adic integers. The section concludes with a statement and sketched proof of the Correspondence Principle, one of the main findings of my earliest work [12, 14]. This relates the integral values taken by $X_{H}$ over $\mathbb{Z}_{p}$ to the periodic and/or divergent points of $H$.
Section 7 begins with a brief review of the $p$-adic Haar probability measure and its use in $p$-adic integration. Several basic $p$-adic integration formulae are introduced, along with the elementary theory of Pontryagin Duality from abstract harmonic analysis and its application to $p$-adic Fourier analysis. Novel content in this section includes Definition 7.3 on page 7.3, which gives the $\ell$-adic characteristic function of $X_{H}$, for varying places $\ell$ of the underlying global field $K$. The section also gives the definition of the $\ell$-adic Wiener algebras. These are the spaces of all functions $K_{\ell}\rightarrow\mathbb{C}$ expressible as the inverse Fourier transform of an $L^{1}$ function $K_{\ell}\rightarrow\mathbb{C}$, and provide a natural functional-analytic backdrop for giving a unified treatment of Fourier analysis across the various completions of $K$. Using the Wiener algebras, we can realize the $\ell$-adic characteristic function of $H$ as the Fourier transform of a self-similar Borel probability measure $d\mu_{H,\ell}$. (Equivalently, this measure is the probability distribution associated to the $\ell$-adic valued random variable induced by $X_{H}$.) The section concludes with some useful formulae for working with $X_{H}$ as an $\ell$-adic valued random variable in the case where $\ell$ is a non-trivial non-archimedean place of $K$.
@@ 3. Global Fields, Absolute Values, & Completions
Proposition 3.1 ($\mu u$ Factorization (Non-Archimedean)).
@@ 5. Hydra Maps
Now we can give the definition of a Hydra map. We begin with the definition of a pre-Hydra.
Definition 5.2.
@@
Remark 5.2.
By linear algebra, we can always identify $K$ with $\mathbb{F}^{d}$ by choosing an $\mathbb{F}$-basis of $K$ and writing $K$’s elements in coordinates. Because of this, everything done here can also be formulated over $\mathbb{F}^{d}$. In that particular case, absolute values on $K$ then become norms on $\mathbb{F}^{d}$; places of $K$ then become equivalence classes of norms on $\mathbb{F}^{d}$.
Remark 5.3.
Though we only consider affine linear maps, the formalism extends just as well to polynomial maps, rational maps, and even analytic maps. I intend to include these in the formulation in a later version of this paper.
Definition 5.3.
@@ A $\Lambda$-Hydra (map) on $K$ is a map $H:\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}$ defined by:
Remark 5.6.
From the perspective of dynamical systems, the integrality property of a Hydra is, by far, its most important property, and plays a critical role in the proof of the Correspondence Principle (Theorem 6.1, given at the end of Section 6).
@@ 6. The Numen of a Hydra Map
Throughout this section, let $H:\mathcal{O}_{K}\rightarrow\mathcal{O}_{K}$ be a $\Lambda$-Hydra map, not necessarily proper, integral, or centered. Following [12, 14], we construct $X_{H}$, the numen of $H$, like so.
Definition 6.1.
@@
Proposition 6.3.
If $H$ is centered, then $X_{H}\left(\mathbf{i}\right)=X_{H}\left(\mathbf{j}\right)$ for all $\mathbf{i}$ and $\mathbf{j}$ belonging to the same equivalence class in $\Sigma_{p}^{*}/\sim$. Hence, $X_{H}:\Sigma_{p}^{*}/\sim\rightarrow K$ is well-defined. We can then view $X_{H}$ as a function $\mathbb{N}_{0}\rightarrow K$ by using $\textrm{DigSum}_{p}$ to identify $\Sigma_{p}^{*}/\sim$ with $\mathbb{N}_{0}$.
Remark 6.2.
If $H$ is not centered, one can still formally define $X_{H}$ without reference to strings by using the functional equations given in Proposition 6.4, below, provided the functional equations admit a solution.
@@
Proposition 6.4 (Siegel (2022); see [12]).
The set of functions $X_{H}:\mathbb{N}_{0}\rightarrow K$ satisfying the functional equations:
@@
(where $r_{j}X_{H}\left(n\right)$ is the image of $X_{H}\left(n\right)\in K$ under $r_{j}\in\textrm{End}_{\mathbb{F}}^{\times}\left(K\right)$) is in a bijective correspondence with the set of $z\in K$ for which $\left(1-r_{0}\right)z=c_{0}$. In particular:
I. If $H$ is proper, then (6.10) has a unique solution, satisfying $X_{H}\left(0\right)=\left(1-r_{0}\right)^{-1}c_{0}$.
II. If $H$ is non-proper, then (6.10) has a solution $X_{H}$ if and only if there are $z\in K$ for which $\left(1-r_{0}\right)z=c_{0}$. For each such $z$, one can set $X_{H}\left(0\right)=z$ and then obtain a distinct solution of (6.10). In this way, we make think of $X_{H}\left(0\right)=z$ as an initial condition subject to which (6.10) can be solved.
@@
In [12], it was shown that, given mild assumptions on the $r_{j}$s, one can use the identification of $\Sigma_{p}^{\infty}/\sim$ with $\mathbb{Z}_{p}$ to interpolate $X_{H}:\mathbb{N}_{0}\rightarrow K$ to a function on $\mathbb{Z}_{p}$. The details for this are given in Lemma 6.1, below. The essential idea is that by making a judicious choice of absolute values on $K$, one can realize $X_{H}$ as a function (or, in general, as a distribution) out of $\mathbb{Z}_{p}$ whose regularity depends on how close the $r_{j}$s are to behaving as contractions with respect to $K$. The behavior of the $r_{j}$s is captured by their norm with respect to a given absolute value of $K$. The result given below is a slightly modified version of the versions originally stated in [14]. The reader should be aware that Lemma 6.1 treats $\mathbb{Z}_{p}$ as a measure space equipped with its real-valued Haar probability measure, and treat $K_{\ell}$ as a measure space equipped with its real-valued Haar measure. We also use the concept of a locally constant function. Readers unfamiliar with either of these can find an expository treatment given of them at the start of Section 7.
Lemma 6.1.
Suppose $H$ is proper. For each non-trivial place $\ell$ of $K$, let:
@@
Now, consider the extension $X_{H}:\mathbb{Z}_{p}\rightarrow K_{\ell}$ defined by the formula:
@@
Remark 6.5.
Like with Proposition 6.4, if $H$ is not proper but there exist $z\in K$ for which $\left(1-r_{0}\right)z=c_{0}$, one can recover the result of Lemma 6.1 subject to the condition that the uniqueness of $X_{H}$ is requires not only the condition (6.21) and the functional equations (6.24), but also that $X_{H}$ satisfy the initial condition $X_{H}\left(0\right)=z$.
@@
which is (6.25), and which establishes the $\ell$-adic boundedness of $X_{H}$.
@@
Theorem 6.1 (Correspondence Principle).
@@ 7. $p$-Adic Fourier Analysis and the $\ell$-adic Characteristic Function of $X_{H}$
Proposition 7.1 (Decomposition formula).
@@
By Lemma 6.1, the condition $\rho_{H,\ell}<1$ guarantees that $X_{H}:\mathbb{Z}_{p}\rightarrow K_{\ell}$ is measurable. As such, we can view $X_{H}$ as a $K_{\ell}$-valued random variable whose associated sample space is the Borel $\sigma$-algebra of $\mathbb{Z}_{p}$. Given any measurable subset $A$ of $K_{\ell}$, the probability $\mathbb{F}\left(X_{H}\in A\right)$ would then be the Haar probability measure of the subset of $\mathbb{Z}_{p}$ on which $X_{H}\in A$.
@@
Proof: By (6.25) from Lemma 6.1, the hypotheses on the $\ell$-adic norms of the $r_{j}$s implies that $X_{H}$ is $\ell$-adically bounded by:
@@
By character orthogonality, using (7.4), we have:
@@
for all $\mathfrak{w}\in\pi_{\ell}^{-B_{H,\ell}}\mathcal{O}_{K_{\ell}}$. Letting $n\geq B_{H,\ell}$, we can apply Proposition 7.5 to get: